Find the solution to each system by the addition (elimination) method.
4s-5t=18
6s+10t=-8

Answer by checkley77(12844)

4s-5t=18 MULTIPLY BY 2 & ADD
6s+10t=-8
8s-10t=36
---------------
14s=28
s=28/14
s=2 ans.
4*2-5t=18
8-5t=18
-5t=18-8
-5t=10
t=10/-5
t=-2 ans.
Proof:
6*2+10*-2=-8
12-20=-8
-8=-8